# Quantum amplification and simulation of strong and ultrastrong coupling of light and matter

### v1

@@ 1.5 The importance of quantum simulations
To overcome significant experimental problems of reaching and coherently controlling the USC and DSC regimes, various quantum simulation methods have been developed Lloyd (1996); Buluta and Nori (2009); Georgescu et al. (2014); Altman et al. (2021). These methods use an easy-to-control quantum system to simulate the properties of a more complex quantum model of interest. More specifically, quantum simulations refer to using controllable quantum systems, such as quantum computers or specialized quantum simulators, to model and understand the behavior of complex quantum systems that are difficult to study using classical computers or analytical methods. These simulations aim to simulate and mimic the behavior of quantum systems in order to gain insights into their properties, dynamics, and interactions.
@@
In this review, we focus on increasing light-matter interactions via quantum simulations, which covers (at least partially) all the above-mentioned applications. As such, we mainly cover methods for simulating the quantum Rabi model in the USC regime by applying drives to the JC model operating in the strong coupling regime. But it is worth noting that the standard and generalized quantum Rabi models enable further simulating and testing large classes of phenomena or even other theories. These include simulating: deterministic quantum nonlinear optics without real photons, but only with virtual photons and single atoms Kockum et al. (2017b, a); Stassi et al. (2017), supersymmetry (SUSY) Tomka et al. (2015), unconventional (polariton-enhanced) superconductivity Sentef et al. (2018), the Higgs mechanism Garziano et al. (2014), and other effects and theories (for a review, see Ref. Kockum et al. (2019)).
@@ 1.6.1 Light-matter interactions amplified by classical drives
As demonstrated theoretically in Ref. Ballester et al. (2012), the quantum Rabi model in the USC regime can be fully simulated by applying two-tone classical drives to the JC model. That method has been demonstrated experimentally in Ref. Lv et al. (2018) by driving a trapped ion by a pair of counter-propagating Raman laser beams. A similar quantum-simulation method has been experimentally implemented with a superconducting qubit embedded in a cavity-QED setup (a microstrip resonator) and driven by two microwave tones Braumüller et al. (2017).
The simulation of the quantum Rabi model by applying strong classical drives (instead of quantum ones) to a JC system enables enhanced-fidelity ultrafast geometric quantum computation Chen et al. (2022). Although a single classical drive applied to a JC system cannot simulate a full quantum Rabi model, it is enough to induce numerous effects, which are usually attributed to the USC regime Wang et al. (2017); Sánchez Muñoz et al. (2020). These include the examples described in detail in Appendix D, i.e., frequency conversion, simultaneous emission of two photons by a single atom, and an analogous effect of the simultaneous excitation of two atoms by a single photon.
@@ 1.6.2 Light-matter interactions amplified by quantum drives
Parametric amplification and quadrature squeezing have a significant role in various fields, including quantum optics Loudon (2000); Scully and Zubairy (1997); Agarwal (2013); Carmichael (2010); Peřina (1991)), atom optics Meystre (2001), and even nonrelativistic and relativistic quantum field theories Itzykson and Zuber (1980). Parametric amplification offers ways to manipulate the quantum properties of light for various purposes, including improving signal-to-noise ratios, enhancing measurement sensitivity, and enabling advanced quantum technologies. We note that the pioneering work of Kennard Kennard (1927) on nonclassical states (which are now referred to as squeezed states) was published almost a century ago, while the first applications of squeezing in quantum metrology, i.e., for gravitational-wave detectors and interferometers were developed over 40 years ago in Refs. Hollenhorst (1979); Caves et al. (1980); Dodonov et al. (1980); Caves (1981). Those works can be considered as the beginning of quantum metrology. To date, the most visible applications of quadrature squeezing are for (i) quantum metrology with squeezed vacuum in the Laser Interferometer Gravitational-Wave Observatory (LIGO) Aasi et al. (2013); Grote et al. (2013); Tse et al. (2019) and the Advanced Virgo Detector Acernese et al. (2019), and (ii) quantum-optical information processing, e.g., in experimental demonstrations of quantum advantage via boson sampling with squeezed light Zhong et al. (2020, 2021); Madsen et al. (2022).
@@
Various applications of the USC regime simulated by applying parametric amplification as a quantum drive have been proposed. For example, giant entangled cat states can be generated in a time-dependent quantum Rabi model simulated by applying parametric amplification to a JC system Chen et al. (2021). Parametric amplification can further enable creating and stabilizing long-lived macroscopic quantum superposition states, not only in a single atom, but also in atomic ensembles Qin et al. (2021), and can even enable ensemble qubits for QIP Qin et al. (2023). It can also be used for beating the so-called 3-dB limit for intracavity squeezing via quantum reservoir engineering Qin et al. (2022). Other examples can be found in the main part of the review.
@@
Squeezing-enhanced interactions between a boson field and matter are not limited to squeezed optical photons. Actually, the first experimental demonstrations of such amplification schemes were realized with squeezed phonons in Boulder Burd et al. (2021, 2023) (see also Ref. Burd et al. (2019)) and squeezed microwave photons in Paris Villiers et al. (2022). More specifically, the Boulder group reported in Ref. Burd et al. (2021) the 3.25-fold amplification of a phonon-mediated interaction between two trapped-ion ${}^{25}$${}^{+}$ hyperfine qubits by parametric modulation of the trapping potential. Another Boulder experiment Burd et al. (2023) showed the interaction between the motional and spin states of a single trapped ${}^{25}$${}^{+}$ ion amplified by phonon squeezing. Phonons in those experimental implementations correspond to a normal mode of trapped-ion motion. The experiments reported in Refs. Burd et al. (2019, 2023) are based on sequential and multiple application of proper squeezing and displacement operations in a close analogy to the theoretical proposal for amplifying Kerr interactions Bartkowiak et al. (2014). Moreover, the Paris-group experiment Villiers et al. (2022) demonstrated two-fold amplified interactions via microwave-photon squeezing (at 5.5 dB) in a superconducting circuit. Specifically, the amplified interaction was observed between a coplanar waveguide resonator capacitively coupled to a transmon qubit. Moreover, a recent theoretical proposal applies the same idea in magnonics, i.e., for amplifying phonon-mediated magnon-spin interactions via virtually-excited squeezed phonons Wang et al. (2023); Nori (2023).
@@ 1.7 Outline of the review
The review is focused on describing USC between photons or phonons and qubits, which are realized by real atoms (like trapped ions) or artificial atoms (e.g., superconducting quantum circuits). Nevertheless, many methods reviewed in this paper can be readily generalized to achieve or simulate the USC between other types of quantum excitations, e.g., between phonons and magnons (see, e.g., Ref. Wang et al. (2023); Nori (2023)), photons and magnons Golovchanskiy et al. (2021), or photons and plasmons Mueller et al. (2020). Thus, the reviewed methods offer new opportunities for quantum technologies also in other fields, like microwave superconducting spintronics or microwave plasmonics.
@@ 2.1 Amplified photon-mechanical interactions in cavity optomechanics
Cavity optomechanics explores the interaction between electromagnetic radiation and mechanical motion Aspelmeyer et al. (2014); Liu et al. (2018). This optomechanical interaction fundamentally originates from the momentum transfer of cavity photons to mechanical objects, referred to as radiation-pressure forces, and can be described by the Hamiltonian Law (1995)
@@
where $g_{0}=Gx_{\rm zpf}$ is the single-photon optomechanical-coupling strength. Usually, $g_{0}$ is extremely weak and, as a result, the ratio $g_{0}/\omega_{m}$ is very small. For these reasons, it is an experimental challenge to observe effects of the detuned optomechanical interaction $H_{\rm OMI}$. Therefore, a large number of methods have been proposed to amplify optomechanical interaction. Below, we review such methods.
@@ 2.1.1 Amplification via linearization
$\dot{b}=\;$ $-i\omega_{m}\omega_{m}+ig_{0}a^{\dagger}a-\frac{\gamma_{m}}{2}b-\sqrt{\gamma_{m}}b_{\rm in}\left(t\right),$ (8)
where $\Delta=\omega_{\rm cav}-\omega_{d}$ is the detuning of the cavity resonance $\omega_{\rm cav}$ from the strong driving field, $\kappa$ is the cavity decay rate, and $\gamma_{m}$ is the mechanical decay rate. Moreover, $a_{\rm in}\left(t\right)$ and $b_{\rm in}\left(t\right)$ are the input-noise annihilation operators for the cavity field and the mechanical motion, respectively.
@@
$\delta\dot{b}=\;$ $-i\omega_{m}\omega_{m}+ig_{0}\left(\alpha\delta a^{\dagger}+\alpha^{*}\delta a\right)-\frac{\gamma_{m}}{2}\delta b-\sqrt{\gamma_{m}}b_{\rm in},$ (16)
@@
is referred to as the linearized optomechanical interaction strength, and $\theta$ is the phase of the amplitude $\alpha$. Here, $\bar{n}_{c}=\left|\alpha\right|^{2}$ is the average number of intracavity photons. Thus $g_{c}$ is enhanced by a factor of $\sqrt{\bar{n}_{c}}$, compared to the bare single-photon coupling $g_{0}$.
@@
In the displaced frame, $H_{-}$ acts as a beam-splitter-like interaction and results in the exchange of a single excitation between the cavity and the mechanical resonator. In the original frame, this exchange corresponds to an anti-Stokes scattering process where a single photon of the driving field is scattered into the cavity resonance, while simultaneously absorbing a mechanical phonon. The Hamiltonian $H_{-}$ has already been widely used for, e.g., sideband cooling of mechanical motion Marquardt et al. (2007); Wilson-Rae et al. (2007); Teufel et al. (2011); Asjad et al. (2016); Clark et al. (2017) and coherent state transfer between the cavity and mechanical modes Fiore et al. (2011); Verhagen et al. (2012); Dong et al. (2012).
@@
In the displaced frame, $H_{+}$ acts as a two-mode-squeezing-like interaction and results in a simultaneous excitation of a cavity photon and a mechanical phonon. In the original frame, this simultaneous excitation corresponds to a Stokes scattering process, where a single photon of the driving field is scattered into the cavity resonance, while simultaneously exciting a mechanical phonon. The Hamiltonian $H_{+}$ has already been widely used for, e.g., quantum-limited amplification Massel et al. (2011); Nunnenkamp et al. (2014) and generating entanglement between the cavity and mechanical modes Vivoli et al. (2016); Hofer et al. (2016); Marinković et al. (2018).
@@ 2.1.2 Resonant amplification
A possible approach is to linearly couple the optomechanical cavity to another empty cavity Stannigel et al. (2012); Ludwig et al. (2012); Kómár et al. (2013), as shown in Fig. 1. Such a double-cavity setup can be realized with, e.g., a membrane-in-the-middle optomechanical system Thompson et al. (2008) or two coupled whispering-gallery-mode microresonators Grudinin et al. (2010); Jing et al. (2014); Zhang et al. (2018a).
Figure 1: (a) Schematics of the double-cavity approach to enhance the single-photon optomechanical interaction. The two cavities, both with frequency $\omega_{\rm cav}$, are coupled through a beam-splitter interaction of strength $J$, which in turn leads to two normal modes with resonance frequencies $\omega_{\pm}=\omega_{\rm cav}\pm J$. The rightmost mirror is oscillating with frequency $\omega_{m}$. (b) Resonant two-mode optomechanical interaction described by $H_{\rm OMI}^{\rm DC}$ in Eq. (24). Under the resonance condition $\omega_{m}\approx 2J$, a photon can be scattered, through the optomechanical interaction $g_{0}$, from the mode $a_{-}$ ($a_{+}$) into the mode $a_{+}$ ($a_{-}$) by the absorption (emission) of a mechanical phonon in the mode $b$. Panels (a) and (b) are reproduced with permission from Ref. Kómár et al. (2013), P. Kómár et al., Phys. Rev. A 87, 013839 (2013).
@@
where $c$ is the annihilation operator for the empty cavity and $J$ is the coupling strength. In the case when these two cavities have the same frequency $\omega_{\rm cav}$, the coupling $H_{\rm IC}$ leads to the formation of the normal modes (often referred to as supermodes)
@@
As demonstrated in Refs. Liao et al. (2014); Liao and Tian (2016), if the coupling strength $J$ in Eq. (21) is assumed to be modulated sinusoidally so that
$J\mapsto J\left(t\right)=\xi\omega_{0}\cos\left(\omega_{0}t\right),$ (28)
@@
$g_{\rm M}=\frac{1}{2}g_{0}J_{2n_{0}}\left(2\xi\right)$ (30)
being an effective optomechanical interaction. Here, $\xi$ is the dimensionless modulation amplitude, $\omega_{0}$ is the modulation frequency, $\delta=\omega_{m}-2n_{0}\omega_{0}$ is the modulation-induced detuning, and $J_{n}\left(z\right)$ is the $n$th-order Bessel function of the first kind. The number $n_{0}$ is a special integer such that the coupling with a detuning $\delta$ is nearly resonant. The superscript “M” refers to the modulation of the intercavity coupling.
The Hamiltonian $H_{\rm OMI}^{\rm M}$ essentially describes an effective force $\propto\langle c^{\dagger}c-a^{\dagger}a\rangle$ acting on a mechanical resonator with an effective frequency $\delta$. Compared to the natural optomechanical interaction $H_{\rm OMI}$, the near-resonant coupling $H_{\rm OMI}^{\rm M}$ can induce a single-photon mechanical displacement proportional to $g_{0}/\left|\delta\right|$ rather than to $g_{0}/\omega_{m}$. This indicates that, as shown in Fig. 2(a), the optomechanical nonlinearity is strongly amplified. The physical reason for this amplification is that the single-photon hopping between the two cavities at the proper times accumulates the displacement effect when the driving force is in phase with the mechanical oscillation. A similar enhancement of the nonlinearity can also be obtained by modulating the cavity frequencies Liao et al. (2015). The resonant amplification based on modulation has been studied for the generation of mechanical Schrödinger cat states via flipping a qubit repeatedly Tian (2005), and even for the detection of the virtual radiation pressure arising from atom-field USC Cirio et al. (2017).
Figure 2: (a) Maximum mechanical displacement, $2g/\left|\delta\right|$, in a double-cavity optomechanical system with a periodically modulated intercavity coupling. Inset: the second-order Bessel function of the first kind, $J_{2}\left(2\xi\right)$, plotted as a function of the modulation amplitude $\xi$. (b) Modified cavity density of states by the residual optomechanical interaction [described by $H_{\delta}$ in Eq. (31)] for the “$+$” polariton mode. The inset shows the asymmetrical density of states for the “--” (left) and “$+$” (right) polariton modes. The solid curves are analytical results; the dashed curve is obtained from a numerical simulation of the master equation. Panels (a) and (b) are adapted with permission, respectively, from Ref. Liao and Tian (2016), J.-Q. Liao et al., Phys. Rev. Lett. 116, 163602 (2016) and Ref. Lemonde et al. (2013), M.-A. Lemonde et al., Phys. Rev. Lett. 111, 053602 (2013).
@@
is neglected. This coupling is of the same form as the standard optomechanical interaction and, as suggested in Refs. Lemonde et al. (2013); Børkje et al. (2013); Kronwald and Marquardt (2013), can be amplified when the coupling $H_{\delta}$ is treated as a perturbation to the linearized Hamiltonian $H_{\rm OMI}^{\rm L}$ in Eq. (17). In the absence of the coupling $H_{\delta}$, the Hamiltonian $H_{\rm OMI}^{\rm L}$ can be diagonalized, yielding

### v2

@@ 1.5 The importance of quantum simulations
To overcome significant experimental problems of reaching and coherently controlling the USC and DSC regimes, various quantum simulation methods have been developed [176, 177, 178, 179, 180]. These methods use an easy-to-control quantum system to simulate the properties of a more complex quantum model of interest. More specifically, quantum simulations refer to using controllable quantum systems, such as quantum computers or specialized quantum simulators, to model and understand the behavior of complex quantum systems that are difficult to study using classical computers or analytical methods. These simulations aim to simulate and mimic the behavior of quantum systems in order to gain insights into their properties, dynamics, and interactions.
@@
In this review, we focus on increasing light-matter interactions via quantum simulations, which covers (at least partially) all the above-mentioned applications. As such, we mainly cover methods for simulating the quantum Rabi model in the USC regime by applying drives to the JC model operating in the strong coupling regime. But it is worth noting that the standard and generalized quantum Rabi models enable further simulating and testing large classes of phenomena or even other theories. These include simulating: deterministic quantum nonlinear optics without real photons, but only with virtual photons and single atoms [163, 162, 85], supersymmetry (SUSY) [181], unconventional (polariton-enhanced) superconductivity [100], the Higgs mechanism [168], and other effects and theories (for a review, see Ref. [49]).
@@ 1.6.1 Light-matter interactions amplified by classical drives
As demonstrated theoretically in Ref. [182], the quantum Rabi model in the USC regime can be fully simulated by applying two-tone classical drives to the JC model. That method has been demonstrated experimentally in Ref. [183] by driving a trapped ion by a pair of counter-propagating Raman laser beams. A similar quantum-simulation method has been experimentally implemented with a superconducting qubit embedded in a cavity-QED setup (a microstrip resonator) and driven by two microwave tones [184].
The simulation of the quantum Rabi model by applying strong classical drives (instead of quantum ones) to a JC system enables enhanced-fidelity ultrafast geometric quantum computation [185]. Although a single classical drive applied to a JC system cannot simulate a full quantum Rabi model, it is enough to induce numerous effects, which are usually attributed to the USC regime [83, 186]. These include the examples described in detail in Appendix D, i.e., frequency conversion, simultaneous emission of two photons by a single atom, and an analogous effect of the simultaneous excitation of two atoms by a single photon.
@@ 1.6.2 Light-matter interactions amplified by quantum drives
Parametric amplification and quadrature squeezing have a significant role in various fields, including quantum optics [187, 1, 2, 188, 3]), atom optics [15], and even nonrelativistic and relativistic quantum field theories [189]. Parametric amplification offers ways to manipulate the quantum properties of light for various purposes, including improving signal-to-noise ratios, enhancing measurement sensitivity, and enabling advanced quantum technologies. We note that the pioneering work of Kennard [190] on nonclassical states (which are now referred to as squeezed states) was published almost a century ago, while the first applications of squeezing in quantum metrology, i.e., for gravitational-wave detectors and interferometers were developed over 40 years ago in Refs. [191, 192, 193, 194]. Those works can be considered as the beginning of quantum metrology. To date, the most visible applications of quadrature squeezing are for (i) quantum metrology with squeezed vacuum in the Laser Interferometer Gravitational-Wave Observatory (LIGO) [195, 196, 197] and the Advanced Virgo Detector [198], and (ii) quantum-optical information processing, e.g., in experimental demonstrations of quantum advantage via boson sampling with squeezed light [199, 200, 201].
@@
Various applications of the USC regime simulated by applying parametric amplification as a quantum drive have been proposed. For example, giant entangled cat states can be generated in a time-dependent quantum Rabi model simulated by applying parametric amplification to a JC system [202]. Parametric amplification can further enable creating and stabilizing long-lived macroscopic quantum superposition states, not only in a single atom, but also in atomic ensembles [203], and can even enable ensemble qubits for QIP [204]. It can also be used for beating the so-called 3-dB limit for intracavity squeezing via quantum reservoir engineering [205]. Other examples can be found in the main part of the review.
@@
Squeezing-enhanced interactions between a boson field and matter are not limited to squeezed optical photons. Actually, the first experimental demonstrations of such amplification schemes were realized with squeezed phonons in Boulder [206, 207] (see also Ref. [208]) and squeezed microwave photons in Paris [209]. More specifically, the Boulder group reported in Ref. [206] the 3.25-fold amplification of a phonon-mediated interaction between two trapped-ion 25Mg+ hyperfine qubits by parametric modulation of the trapping potential. Another Boulder experiment [207] showed the interaction between the motional and spin states of a single trapped 25Mg+ ion amplified by phonon squeezing. Phonons in those experimental implementations correspond to a normal mode of trapped-ion motion. The experiments reported in Refs. [208, 207] are based on sequential and multiple application of proper squeezing and displacement operations in a close analogy to the theoretical proposal for amplifying Kerr interactions [210]. Moreover, the Paris-group experiment [209] demonstrated two-fold amplified interactions via microwave-photon squeezing (at 5.5 dB) in a superconducting circuit. Specifically, the amplified interaction was observed between a coplanar waveguide resonator capacitively coupled to a transmon qubit. Moreover, a recent theoretical proposal applies the same idea in magnonics, i.e., for amplifying phonon-mediated magnon-spin interactions via virtually-excited squeezed phonons [211, 212].
@@ 1.7 Outline of the review
The review is focused on describing USC between photons or phonons and qubits, which are realized by real atoms (like trapped ions) or artificial atoms (e.g., superconducting quantum circuits). Nevertheless, many methods reviewed in this paper can be readily generalized to achieve or simulate the USC between other types of quantum excitations, e.g., between phonons and magnons (see, e.g., Ref. [211, 212]), photons and magnons [213], or photons and plasmons [161]. Thus, the reviewed methods offer new opportunities for quantum technologies also in other fields, like microwave superconducting spintronics or microwave plasmonics.
@@ 2.1 Amplified photon-mechanical interactions in cavity optomechanics
Cavity optomechanics explores the interaction between electromagnetic radiation and mechanical motion [214, 215]. This optomechanical interaction fundamentally originates from the momentum transfer of cavity photons to mechanical objects, referred to as radiation-pressure forces, and can be described by the Hamiltonian (hereafter we set $\hbar=1$[216]
@@
where $g_{0}=Gx_{\rm zpf}$ is the single-photon optomechanical-coupling strength. A necessary condition for Eq. (5) to be valid is that $g_{0}$ is much smaller than $2\omega_{\rm cav}\pm\omega_{m}$, such that the mechanically induced creation and annihilation of photon pairs can be neglected. In most experimental situations, this condition is well satisfied. Indeed, $g_{0}$ is extremely weak and, as a result, the ratio $g_{0}/\omega_{m}$ is very small. For these reasons, it is an experimental challenge to observe effects of the detuned optomechanical interaction $H_{\rm OMI}$. Therefore, a large number of methods have been proposed to amplify the optomechanical interaction. Below, we review such methods.
@@ 2.1.1 Amplification via linearization
$\dot{b}=\;$ $-i\omega_{m}b+ig_{0}a^{\dagger}a-\frac{\gamma_{m}}{2}b-\sqrt{\gamma_{m}}b_{\rm in}\left(t\right),$ (8)
where $\Delta=\omega_{\rm cav}-\omega_{d}$ is the detuning of the cavity resonance $\omega_{\rm cav}$ from the strong driving field, $\kappa$ is the cavity decay rate, and $\gamma_{m}$ is the mechanical decay rate. Moreover, $a_{\rm in}\left(t\right)$ and $b_{\rm in}\left(t\right)$ are the input-noise annihilation operators for the cavity field and the mechanical motion, respectively. Note that here, to derive the quantum Langevin equations of motion in Eqs. (7) and (8), the Weisskopf-Wigner approximation has been made [1], such that the decay rates $\kappa$ and $\gamma_{m}$ are constant.
@@
$\delta\dot{b}=\;$ $-i\omega_{m}\delta b+ig_{0}\left(\alpha\delta a^{\dagger}+\alpha^{*}\delta a\right)-\frac{\gamma_{m}}{2}\delta b-\sqrt{\gamma_{m}}b_{\rm in},$ (16)
@@
is referred to as the linearized optomechanical interaction strength, and $\theta$ is the phase of the amplitude $\alpha$. Here, $\bar{n}_{c}=\left|\alpha\right|^{2}$ is the average number of intracavity photons. Thus $g_{c}$ is enhanced by a factor of $\sqrt{\bar{n}_{c}}$, compared to the bare single-photon coupling $g_{0}$. It has been experimentally shown that such an enhanced coupling can even reach the regime of the USC [217].
@@
In the displaced frame, $H_{-}$ acts as a beam-splitter-like interaction and results in the exchange of a single excitation between the cavity and the mechanical resonator. In the original frame, this exchange corresponds to an anti-Stokes scattering process where a single photon of the driving field is scattered into the cavity resonance, while simultaneously absorbing a mechanical phonon. The Hamiltonian $H_{-}$ has already been widely used for, e.g., sideband cooling of mechanical motion [218, 219, 220, 221, 222] and coherent state transfer between the cavity and mechanical modes [223, 224, 225].
@@
In the displaced frame, $H_{+}$ acts as a two-mode-squeezing-like interaction and results in a simultaneous excitation of a cavity photon and a mechanical phonon. In the original frame, this simultaneous excitation corresponds to a Stokes scattering process, where a single photon of the driving field is scattered into the cavity resonance, while simultaneously exciting a mechanical phonon. The Hamiltonian $H_{+}$ has already been widely used for, e.g., quantum-limited amplification [226, 227] and generating entanglement between the cavity and mechanical modes [228, 229, 230].
@@ 2.1.2 Resonant amplification
A possible approach is to linearly couple the optomechanical cavity to another empty cavity [231, 232, 233], as shown in Fig. 1. Such a double-cavity setup can be realized with, e.g., a membrane-in-the-middle optomechanical system [234] or two coupled whispering-gallery-mode microresonators [235, 236, 237].
Figure 1: (a) Schematics of the double-cavity approach to enhance the single-photon optomechanical interaction. The two cavities, both with frequency $\omega_{\rm cav}$, are coupled through a beam-splitter interaction of strength $J$, which in turn leads to two normal modes with resonance frequencies $\omega_{\pm}=\omega_{\rm cav}\pm J$. The rightmost mirror is oscillating with frequency $\omega_{m}$. (b) Resonant two-mode optomechanical interaction described by $H_{\rm OMI}^{\rm DC}$ in Eq. (24). Under the resonance condition $\omega_{m}\approx 2J$, a photon can be scattered, through the optomechanical interaction $g_{0}$, from the mode $a_{-}$ ($a_{+}$) into the mode $a_{+}$ ($a_{-}$) by the absorption (emission) of a mechanical phonon in the mode $b$. Panels (a) and (b) are reproduced with permission from Ref. [233], P. Kómár et al., Phys. Rev. A 87, 013839 (2013).
@@
where $c$ is the annihilation operator for the empty cavity and $J$ is the coupling strength. Note that the presence of the negative sign does not affect the results of interest. In the case when these two cavities have the same frequency $\omega_{\rm cav}$, the coupling $H_{\rm IC}$ leads to the formation of the normal modes (often referred to as supermodes)
@@
As demonstrated in Refs. [238, 239], if the coupling strength $J$ in Eq. (21) is assumed to be modulated sinusoidally so that
$J\mapsto J\left(t\right)=\zeta\omega_{0}\cos\left(\omega_{0}t\right),$ (28)
@@
$g_{\rm M}=\frac{1}{2}g_{0}J_{2n_{0}}\left(2\zeta\right)$ (30)
being an effective optomechanical interaction. Here, $\zeta$ is the dimensionless modulation amplitude, $\omega_{0}$ is the modulation frequency, $\delta=\omega_{m}-2n_{0}\omega_{0}$ is the modulation-induced detuning, and $J_{n}\left(z\right)$ is the $n$th-order Bessel function of the first kind. The number $n_{0}$ is a special integer such that the coupling with a detuning $\delta$ is nearly resonant. The superscript “M” refers to the modulation of the intercavity coupling.
The Hamiltonian $H_{\rm OMI}^{\rm M}$ essentially describes an effective force $\propto\langle c^{\dagger}c-a^{\dagger}a\rangle$ acting on a mechanical resonator with an effective frequency $\delta$. Compared to the natural optomechanical interaction $H_{\rm OMI}$, the near-resonant coupling $H_{\rm OMI}^{\rm M}$ can induce a single-photon mechanical displacement proportional to $g_{0}/\left|\delta\right|$ rather than to $g_{0}/\omega_{m}$. This indicates that, as shown in Fig. 2(a), the optomechanical nonlinearity is strongly amplified. The physical reason for this amplification is that the single-photon hopping between the two cavities at the proper times accumulates the displacement effect when the driving force is in phase with the mechanical oscillation. A similar enhancement of the nonlinearity can also be obtained by modulating the cavity frequencies [240]. The resonant amplification based on modulation has been studied for the generation of mechanical Schrödinger cat states via flipping a qubit repeatedly [241], and even for the detection of the virtual radiation pressure arising from atom-field USC [242].
Figure 2: (a) Maximum mechanical displacement, $2g/\left|\delta\right|$, in a double-cavity optomechanical system with a periodically modulated intercavity coupling. Inset: the second-order Bessel function of the first kind, $J_{2}\left(2\zeta\right)$, plotted as a function of the modulation amplitude $\zeta$. (b) Modified cavity density of states by the residual optomechanical interaction [described by $H_{\delta}$ in Eq. (31)] for the “$+$” polariton mode. The inset shows the asymmetrical density of states for the “--” (left) and “$+$” (right) polariton modes. The solid curves are analytical results; the dashed curve is obtained from a numerical simulation of the master equation. Panels (a) and (b) are adapted with permission, respectively, from Ref. [239], J.-Q. Liao et al., Phys. Rev. Lett. 116, 163602 (2016) and Ref. [243], M.-A. Lemonde et al., Phys. Rev. Lett. 111, 053602 (2013).
@@
is neglected. This coupling is of the same form as the standard optomechanical interaction and, as suggested in Refs. [243, 244, 245], can be amplified when the coupling $H_{\delta}$ is treated as a perturbation to the linearized Hamiltonian $H_{\rm OMI}^{\rm L}$ in Eq. (17). In the absence of the coupling $H_{\delta}$, the Hamiltonian $H_{\rm OMI}^{\rm L}$ can be diagonalized, yielding